# Convergence Acceleration of Favre-Averaged Non-Linear Harmonic Method

### v1

@@ Abstract
This paper develops a numerical procedure to accelerate the convergence of the Favre-averaged Non-Linear Harmonic (FNLH) method. The scheme provides a unified mathematical framework for solving the sparse linear systems formed by the mean flow and the time-linearized harmonic flows of FNLH in an explicit or implicit fashion. The approach explores the similarity of the sparse linear systems of FNLH and leads to a memory efficient procedure, so that its memory consumption does not depend on the number of harmonics to compute. The proposed method has been implemented in the industrial CFD solver HYDRA. Two test cases are used to conduct a comparative study of explicit and implicit schemes in terms of convergence, computational efficiency, and memory consumption. Comparisons show that the implicit scheme yields better convergence than the explicit scheme and is also roughly 7 to 10 times more computationally efficient than the explicit scheme with 4 levels of multigrid. Furthermore, the implicit scheme consumes only approximately $50\%$ of the explicit scheme with four levels of multigrid. Compared with the full annulus unsteady Reynolds averaged Navier-Stokes (URANS) simulations, the implicit scheme produces comparable results to URANS with computational time and memory consumption that are two orders of magnitude smaller.
@@ Introduction
Computational Fluid Dynamics (CFD) has been routinely used in the aerodynamic design of fluid machinery. A fluid machinery (i.e. turbomachinery) can involve a series of rotational and stationary rows of blades to exchange energy between the machine and the working fluid. Even though the exchange of energy between the working fluid and the machine is dependent on the unsteadiness of the flow [1], steady Reynolds Averaged Navier-Stokes (RANS) has been the workhorse in the industry for designing turbomachinery. However, unsteady interactions among turbomachinery components can be significant and should be considered at the design stage. Unsteady RANS (URANS) simulation is an approach to model these unsteady interactions, but they can increase the computational cost by several orders of magnitude compared to the steady RANS approach. The industry has had a significant interest in developing computationally efficient design methods to model these unsteady interactions. This work is a step in this direction.
@@
If the effect of stochastic flow perturbations on the mean flow can be approximated by a suitable turbulence model, the unsteady flow in a turbomachine can be effectively decomposed into a time-mean flow and an unsteady periodic perturbation. Periodic fluctuations are deterministic and can be computed efficiently by harmonic methods, which transform unsteady time domain problems into steady problems in the frequency domain. This removes the constraint of global time stepping in the nonlinear time marching approach (i.e. URANS). For multistage simulations, phase-lag boundaries can be applied to the periodic boundaries, and this effectively avoids the need to model multiple passages or the whole wheel in the simulations. Nevertheless, it is noted that this approach ignores the interactions between the stochastic and the periodic perturbations; besides, the frequencies of the periodic perturbations normally need to be known a priori. Addressing these limitations is out of the scope of this work and will be investigated in future work.
@@
This work concerns the family of harmonic methods that couples a mean flow solver and a time-linearized NS flow solver in the frequency domain. The coupling between the mean flow and the perturbations can be evaluated using the flow perturbations of the linearized NS solver and the mean flow variables. He and Ning [2] proposed the first type of such approach, which is called the Non-Linear Harmonic (NLH) method, but the authors did not elucidate the conversion among conservative and primitive variables in the mean flow. Wang and di Mare [3] proposed a more elaborate and rigorous approach to include the non-linear coupling between the mean flow and the flow perturbations for compressible flows, and the approach is termed the Favre-averaged NLH (FNLH) method. FNLH can be used to simulate unsteady rotor-stator interactions [3], multi-row [4] interactions and more recently multi-shaft [5] unsteady interactions. Furthermore, FNLH can also be used to efficiently simulate non-uniform flow fields in an assembly of bladerows with non-uniform geometries [6, 7]. The other family of harmonic methods is the harmonic balance method [8]. This approach does not require a linearized NS solver, but it solves a finite set of steady-state solutions at fixed time levels in the time domain simultaneously. All of these time-level solutions are coupled through periodic boundaries and a spectral source term. Readers can refer to Hall et al. [9] for more details on this approach.
@@
Harmonic methods (i.e. FNLH) use pseudo-time stepping to solve the unsteady problem in the Fourier space [3]. For each pseudo-time step, the mean flow and the time linearized flow need to be computed. Convergence acceleration techniques are crucial to the computational performance of harmonic methods. NLH used an explicit scheme with geometric multigrid to accelerate its convergence. To the best knowledge of the authors, there is no previous research in the public domain on developing an implicit scheme for NLH or FNLH. However, it is worth mentioning that there has been previous work on developing implicit schemes to accelerate the convergence of the Harmonic Balance (HB) method, such as the work of Ref [10, 11, 12, 13], as harmonic balance method does not need to solve a dedicated time-linearized solver, and hence these approaches cannot be used directly to accelerate the convergence of FNLH or NLH.
@@
There has been extensive research on the acceleration of convergence of a steady NS flow solver  [14, 15, 16] and also a time-linearized flow solver [17] using explicit or implicit approaches. This work will bridge the gap between convergence acceleration techniques in these two aspects and develop an implicit scheme to accelerate the convergence of FNLH. This is the first contribution of this paper. The second contribution of this paper is to formulate a unified formulation to solve the sparse linear systems of FNLH in an explicit or implicit fashion. The similarity of the sparse pattern for the linear systems of the mean flow and the time-linearized flow is explored to develop a memory-efficient scheme. The third contribution is to conduct a comparative study of explicit and implicit schemes regarding convergence, computational efficiency, and memory usage. The proposed method is implemented in the industrial CFD solver HYDRA [18], which a major aero-engine manufacturer routinely uses for aerothermal design.
@@
The paper is organized as follows: the numerical algorithms of HYDRA are briefly introduced in the first place, and the FNLH method is then described. The generalized RK scheme for FNLH is then detailed. The performance of this generalized explicit/implicit RK scheme is demonstrated. This is followed by conclusions and future work.
@@
The numerics of HYDRA are briefly described here, and readers can refer to Ref [19, 17, 18] for more details. HYDRA is a finite-volume CFD solver and uses the edge-based data structure, and this can be illustrated by Fig. 1. The solver is second-order accurate in space. Inviscid flux is calculated with the Jameson-Schmidt-Turkel (JST) scheme using matrix dissipation [20]. The viscous flux is computed using the central difference. To calculate the flow gradient, the gradient on each edge $(i,j)$ is first calculated using a mixture of the central difference and the average flow gradient at nodes $i$ and $j$. Gradient estimates for each node are given by the Green-Gauss gradients [18]. A set of turbulence models is implemented in HYDRA. In this work, the Spalart-Allmaras (SA) [21] turbulent model with wall function is used. With respect to time integration, the hybrid five-stage Runge-Kutta (RK) scheme is implemented, where the non-linear viscous residuals are evaluated only at odd stages. This scheme has been widely used in the literature and demonstrated good performance [22, 18, 23] . To accelerate the convergence of the non-linear solver, geometric multigrid [24] and incomplete matrix factorization [25] are implemented. The former leads to an explicit RK scheme, and the latter leads to an implicit RK scheme.
HYDRA also has an existing time-linearized flow solver and TAPENADE [26] is used to perform the source code transformation to automatically differentiate the non-linear solver. Only the subroutines that contribute to the right-hand side of the nonlinear solver are differentiated, but the time-stepping for the time-linearized solver is still coded manually. Although a time-linearized flow solver solves the real and imaginary parts of the Fourier components of the flow perturbations, the subroutines of the differentiated code from TAPENDA are for flow variables in real numbers and they will be used sequentially to evaluate the real and imaginary parts of the residuals of the time-linearized solver . The explicit RK scheme with multigrid is used to accelerate the convergence of the time-linearized solver, but no implicit solver is implemented so far. FNLH requires a mean flow solver and a time-linearized flow solver, both of which are available in HYDRA, therefore, implementing FNLH in HYDRA does not require a significant modification of the code.
@@
The Favre-averaged Non-Linear Harmonic (FNLH) method is a computational framework which efficiently couples the mean flow and its finite-amplitude periodic perturbations for compressible flows in the frequency domain. Here we provide a brieft description of FNLH and the details of FNLH can be found in Wang and di Mare [3, 4, 6].
@@
in which $\hat{\mathbf{W}}_{i}$ is the Favre-averaged primitive variable and ${\mathbf{W}}^{\prime\prime}_{i}$ is the flow perturbation relative to the Favre-averaged value. It is noted that, for Favre-averaging, density $\rho$ and pressure $p$ are still time-averaged. Their perturbations are denoted as $\rho^{\prime}$ and $p^{\prime}$. The conversion between Favre-averaged and time-averaged conservative variables $\bar{\mathbf{Q}}_{i}=(\bar{\rho},\overline{\rho u},\overline{\rho v},\overline{\rho w},\overline{\rho E})^{T}$ is trivial [27], for example $\overline{\rho u}=\bar{\rho}\tilde{u}$.
In the context of periodic unsteady flows in turbomachinery, considering the schematic in Fig. 2, which features a stator-rotor-stator setting. The flow field of B3 can be decomposed to the combination of a passage-averaged time-mean value, an unsteady disturbance and a stationary passage-to-passage disturbance. If we consider the stationary disturbance as a ”slowly moving” unsteady disturbance and its frequency is approaching zero. The flow field in B3 can be decomposed as a passage-averaged time-mean value $\tilde{\phi}$ and an unsteady flow disturbance $\phi^{\prime\prime}$, whose frequency can be zero.
@@
If the perturbation of a dummy variable ${\phi}^{\prime\prime}$ of control volume $i$ can be represented as the harmonic form:
$\phi^{\prime\prime}_{i}\approx\sum^{l=N_{h}}_{l=-N_{h}}\hat{\phi}_{i,l}e^{I\omega_{i,l}t}$ (2)
@@
where $l$ is the harmonic index, $N_{h}$ is the number of harmonics, $I$ is the imaginary unit $\sqrt{-1}$, $\omega_{i,l}$ is the angular frequency of the $l^{\text{th}}$ harmonic and $t$ is time. Using the flow decomposition and the representation of the flow perturbation in the harmonic form, the coupled system of mean flow and harmonic flow perturbations can be obtained, and this can be represented symbolically in Equation Favre-averaged Non-Linear Harmonic Method:
$\mathbf{R}_{i}+\text{DF}_{i}$ $=0$
$I\omega_{i,l}\mathbf{M}V_{i}\mathbf{\hat{W}}_{i,l}+\frac{\partial\mathbf{R}_{i}}{\partial\mathbf{W}_{i}}\mathbf{\hat{W}}_{i,l}$ $=0$ (3)
@@
in which $l$ is the harmonic index, $V_{i}$ is the volume of the control volume $i$ and $\mathbf{M}$ is the matrix that converts primitive variable $\mathbf{W}$ to conservative variables $\mathbf{Q}$. $\mathbf{R}_{i}$ is the non-linear residual of the mean flow and $\frac{\partial\mathbf{R}_{i}}{\partial\mathbf{W}_{i}}\mathbf{W}^{\prime\prime}_{i}$ the linearized fluxes. DF is the deterministic fluxes [28]. It represents the coupling between the mean flow and harmonic flow perturbations and can be determined using the mean flow variable and harmonic flow perturbations [3]. It should be noted that HYDRA works with primitive variables, and this is why the coupled system in Equation Favre-averaged Non-Linear Harmonic Method is written in the form of primitive variables, and the formulation of FNLH is favoured in HYDRA.
In a multistage setting of FNLH, bladerow interfaces for the mean flow and the time-linearized harmonic flow are treated separately. For the mean flow, the original mixing plane formulation [29] is corrected by DF to take into account the unsteady effect [3]. For the time-linearized flow, a unified formulation has been developed to transmit disturbances in a multi-row environment [4]. Quasi-3D non-reflective treatment [30] are applied to the mean flow and time-linearized flow.
@@
Regarding the general process of FNLH, for the $n^{th}$ non-linear iteration (or outer iteration), the harmonic flow perturbations are evaluated in the first place, DF is then computed using the available flow perturbations in the harmonic form, and the mean flow is then updated with the inclusion of DF. It is noted that the cross-coupling of the harmonics is ignored in the current work. This will be included in our future work.
Convergence Acceleration Procedure of FNLH
This work devises a unified formulation to accelerate the convergence of FNLH. The scheme allows both explicit and implicit schemes to be implemented. Although the mean flow and the time-linearized harmonic flow are solved separately, it will be shown that they can be connected with each other due to the identical sparse patterns of their linear systems. This feature will also be explored to develop a memory-efficient scheme.
Generalized RK Scheme
@@
HYDRA uses the generalized Runge-Kutta (RK) scheme for time integration, both explicit [24] and implicit [31] time integration are available. However, this procedure is only for steady flows, and in this paper it is extended to solve the mean flow and the time-linearized harmonic flows of FNLH. The generalized RK scheme uses the hybrid five-stage scheme RK (5,3), where the non-linear viscous residuals are evaluated at odd stages. It can be written as:
$\mathbf{Q}^{n,(0)}_{i}$ $=\mathbf{Q}^{n}_{i}$
$\mathbf{Q}^{n,(k)}_{i}$ $=\mathbf{Q}^{n,(0)}_{i}-\Gamma\alpha^{(k)}\mathbf{K}^{-1}(\mathbf{R}^{(k)}_{i}-\mathbf{R}^{*}_{i}),k=[1:5]$
$\mathbf{Q}^{n+1}_{i}$ $=\mathbf{Q}^{n,(5)}_{i}$ (4)
@@
In Equation Generalized RK Scheme, $n$ is the current pseudo-time step index, $k$ is the RK stage index, $\mathbf{R}^{(k)}_{i}$ is the non-linear residual of the $k^{\text{th}}$ RK stage, $\mathbf{R}^{*}_{i}$ is the additional term related to the multigrid operation. Its formulation can be found in Monier [18] and for the implicit scheme, this term is set to zero. $\Gamma$ can be considered as a relaxation factor and $\mathbf{K}$ is the matrix preconditioner. To improve the convergence rate and stability, the non-linear residual of the $k^{\text{th}}$ stage $\mathbf{R}^{(k)}_{i}$ is decomposed into two components, namely the inviscid contribution $\mathbf{R}^{(k),inv}_{i}$ and the dissipative or viscous contribution $\mathbf{R}^{(k),vis}_{i}$:
$\mathbf{R}^{(k)}_{i}$ $=\mathbf{R}^{(k),inv}_{i}+\mathbf{R}^{(k),vis}_{i}$ (5)
@@
$\mathbf{R}^{(k),inv}_{i}$ $=\mathbf{R}^{inv}_{i}(\mathbf{Q}^{n,k-1)})$ (6)
$\mathbf{R}^{(k),vis}_{i}$ $=\beta_{k}\mathbf{R}^{vis}_{i}(\mathbf{Q}^{n,k-1})+(1-\beta_{k})\mathbf{R}^{(k-1),vis}_{i}$ (7)
$\mathbf{R}^{inv}_{i}(\mathbf{Q}^{n,k-1})$ is the inviscid contribution using the flow variable $\mathbf{Q}^{n,k-1}$, $\mathbf{R}^{(k),vis}_{i}$ is the viscous contribution and is a mixture of the viscous flux using the flow variable $\mathbf{Q}^{n,k-1}_{i}$ and the viscous residual $\mathbf{R}^{(k-1),vis}_{i}$ from the previous RK stage. The coefficients $\alpha^{(k)}$ and $\beta^{(k)}$ are summarized in Table 1. These coefficients are optimized to damp high-frequency errors and lead to better convergence when used with multigrid. Furthermore, these coefficients allow the CFL number of the explicit scheme to be increased to 2.

### v2

@@ Abstract
This paper develops a numerical procedure to accelerate the convergence of the Favre-averaged Non-Linear Harmonic (FNLH) method. The scheme provides a unified mathematical framework for solving the sparse linear systems formed by the mean flow and the time-linearized harmonic flows of FNLH in an explicit or implicit fashion. The approach explores the similarity of the sparse linear systems of FNLH and leads to a memory efficient procedure, so that its memory consumption does not depend on the number of harmonics to compute. The proposed method has been implemented in the industrial CFD solver Hydra. Three test cases are used to conduct a comparative study of explicit and implicit schemes in terms of convergence, computational efficiency, and memory consumption. Comparisons show that the implicit scheme yields better convergence than the explicit scheme and is also roughly 7 to 10 times more computationally efficient than the explicit scheme with 4 levels of multigrid. Furthermore, the implicit scheme consumes only approximately $50\%$ of the memory required by the explicit scheme with four levels of multigrid. Compared with the full annulus unsteady Reynolds averaged Navier-Stokes (URANS) simulations, the implicit scheme produces comparable results to URANS with computational time and memory consumption that are two orders of magnitude smaller.
@@ Introduction
Computational Fluid Dynamics (CFD) has been routinely used in the aerodynamic design of fluid machinery. A fluid machinery (i.e. turbomachinery) can involve a series of rotational and stationary rows of blades to exchange energy between the machine and the working fluid. Although the exchange of energy between the working fluid and the machine is dependent on the unsteadiness of the flow [1], steady Reynolds Averaged Navier-Stokes (RANS) has been the workhorse in the industry to design turbomachinery. However, unsteady interactions among turbomachinery components can be significant and should be considered at the design stage. Unsteady RANS (URANS) simulation is an approach to model these unsteady interactions, but they can increase the computational cost by several orders of magnitude compared to the steady RANS approach. The industry has shown significant interest in developing computationally efficient design methods to model these unsteady interactions. This work is a step in this direction.
@@
If the effect of stochastic flow perturbations on the mean flow can be approximated by a suitable turbulence model, the unsteady flow in a turbomachine can be effectively decomposed into a time-mean flow and an unsteady periodic perturbation. Periodic fluctuations are deterministic and can be computed efficiently by harmonic methods, which transform unsteady time domain problems into steady problems in the frequency domain. This removes the constraint of global time stepping in the nonlinear time marching approach (i.e. URANS). For multistage simulations, phase-lag boundaries can be applied to the periodic boundaries, and this effectively avoids the need to model multiple passages or the whole wheel in the simulations. Nevertheless, this approach ignores the interactions between stochastic and periodic perturbations; besides, the frequencies of the periodic perturbations normally need to be known a priori.
@@
This work concerns the family of harmonic methods that couples a mean flow solver and a time-linearized NS flow solver in the frequency domain. The coupling between the mean flow and the perturbations can be evaluated using the flow perturbations of the linearized NS solver and the mean flow variables. He and Ning [2] proposed the first type of such approach, which is called the Non-Linear Harmonic (NLH) method, but the authors did not elucidate the conversion among time-averaged conservative and primitive variables in the mean flow. Wang and di Mare [3] proposed a more elaborate and rigorous approach to include the nonlinear coupling between the mean flow and the flow perturbations for compressible flows, and the approach is termed the Favre-averaged NLH (FNLH) method. FNLH can be used to simulate unsteady rotor-stator interactions [3], multi-row [4] interactions and more recently multi-shaft [5] unsteady interactions. Furthermore, FNLH can also be used to efficiently simulate non-uniform flow fields in an assembly of bladerows with non-uniform geometries [6, 7]. The other family of harmonic methods is the Harmonic Balance (HB) method [8]. This approach does not require a linearized NS solver, but it solves a finite set of steady-state solutions at fixed time levels in the time domain simultaneously. All of these time-level solutions are coupled through periodic boundaries and a spectral source term. Readers can refer to Hall et al. [9] for more details on this approach.
@@
Harmonic methods (i.e., FNLH) use pseudo-time stepping to solve the unsteady problem in the Fourier space [3]. For each pseudo-time step, the mean flow and the time-linearized flow need to be computed. Convergence acceleration techniques are crucial to the computational performance of harmonic methods. NLH used an explicit scheme with geometric multigrid to accelerate its convergence [2]. However, there is no previous research in the public domain on developing an implicit scheme for NLH or FNLH. On the other hand, it is worth mentioning that there has been previous work on developing implicit schemes to accelerate the convergence of the HB method, the fully implicit approach [10, 11] has been proposed but memory usage can increase rapidly as more harmonics are computed. Considering the fact that the coupling among the harmonics and the mean flow can be weak, a segmented approach [12, 13] can be used to solve the mean flow and each harmonic perturbation separately, so that the memory requirement depends weakly on the number of harmonics to compute. However, since the HB method does not require a dedicated time-linearized solver, these approaches cannot be used directly to accelerate the convergence of FNLH or NLH.
@@
There has been extensive research on the acceleration of convergence of a steady NS flow solver  [14, 15, 16] and also a time-linearized flow solver [17] using explicit or implicit approaches. This work will bridge the gap between convergence acceleration techniques in these two aspects and develop an implicit scheme to accelerate the convergence of FNLH. This is the first contribution of this paper. The second contribution of this paper is to formulate a unified formulation to solve the sparse linear systems of FNLH in an explicit or implicit fashion. The third contribution is to conduct a comparative study of explicit and implicit schemes with respect to convergence, computational efficiency, and memory usage. The proposed method is implemented in the industrial CFD solver Hydra [18], which a major aero-engine manufacturer routinely uses for aerothermal design.
@@
The paper is organized as follows: the numerical algorithms of Hydra are briefly introduced in the first place, and the FNLH method is then briefly described. The generalized RK scheme for FNLH is then detailed. The performance of this generalized explicit/implicit RK scheme is demonstrated. This is followed by conclusions and future work.
@@
The numerics of Hydra are briefly described here, and readers can refer to Ref. [19, 17, 18] for more details. Hydra is a finite-volume CFD solver and uses the edge-based data structure, and this can be illustrated by Fig. 1. The solver is second-order accurate in space. Inviscid flux is calculated with the Jameson-Schmidt-Turkel (JST) scheme using matrix dissipation [20]. The viscous flux is computed using the central difference. To calculate the flow gradient, the gradient on each edge $(i,j)$ is first calculated using a mixture of the central difference and the average flow gradient at nodes $i$ and $j$. Gradient estimates for each node are given by the Green-Gauss gradients [18]. In Hydra, a set of turbulence models is implemented. In this work, the Spalart-Allmaras (SA) [21] turbulence model with wall function is used. With respect to time integration, A low-dissipation and low-dispersion RK scheme RK($s,p$) is implemented in Hydra, where $s$ is the number of stages and $p$ is the number of evaluations of the viscous contribution per time step. In Hydra $s=5$ and $p=3$. This scheme has been widely used in the literature and has demonstrated good performance [22, 18, 23] . To accelerate the convergence of the nonlinear solver, the geometric multigrid [24] and the incomplete matrix factorization [25] have been implemented. The former leads to an explicit RK scheme, and the latter leads to an implicit RK scheme.
Hydra has an existing time-linearized harmonic flow solver. TAPENADE [26] is used to perform the source code transformation to automatically differentiate the nonlinear solver and produce the linearized code. Only the subroutines that contribute to the non-linear residuals are differentiated, but the time-stepping for the time-linearized solver is still coded manually and similar to the non-linear solver the RK(5,3) scheme is used. The linearized code can only work with real numbers; therefore, the real and imaginary parts of harmonic flow perturbations are stored separately. With respect to turbulence modeling, although source code transformation can also be applied to subroutines that are related to turbulence models, frozen turbulence is still used in this work for enhanced robustness. Future work is planned to remove this restriction.
In the baseline Hydra, the time-linearized harmonic flow solver is independent from the non-linear solver. Furthermore, it can only compute one harmonic at a time for a single bladerow. As FNLH requires one to compute the mean flow and the harmonic flow simultaneously for multiple harmonics and multiple bladerows, the limitations of the baseline Hydra have been completely removed in the current code development to successfully implement the FNLH method.
@@
The Favre-averaged Non-Linear Harmonic (FNLH) method is a computational framework that efficiently couples the mean flow and its finite-amplitude periodic perturbations for compressible flows in the frequency domain. Here we provide a brief description of FNLH and the details of FNLH can be found in Wang and di Mare [3, 4, 6].
@@
in which $\hat{\mathbf{W}}_{i}$ is the Favre-averaged primitive variable and ${\mathbf{W}}^{\prime\prime}_{i}$ is the flow perturbation relative to the Favre-averaged value. It is noted that, for Favre-averaging, density $\rho$ and pressure $p$ are still time-averaged [27]. Their perturbations are denoted as $\rho^{\prime}$ and $p^{\prime}$. The conversion between Favre-averaged and time-averaged conservative variables $\bar{\mathbf{Q}}_{i}=(\bar{\rho},\overline{\rho u},\overline{\rho v},\overline{\rho w},\overline{\rho E})^{T}$ is trivial [28], for example $\overline{\rho u}=\bar{\rho}\tilde{u}$.
In the context of periodic unsteady flows in turbomachinery, considering the schematic in Fig. 2, which features a stator(B1) - rotor(B2) - stator(B3) setting. The flow field of B3 can be decomposed to the combination of a passage-averaged time-mean value, an unsteady disturbance, and a stationary passage-to-passage disturbance. If we consider the stationary disturbance as a ”slowly moving” unsteady disturbance and its frequency is approaching zero. The flow field in B3 can be decomposed as a passage-averaged time-mean value $\tilde{\phi}$ and an unsteady flow disturbance $\phi^{\prime\prime}$, whose frequency can be zero. Here, $\phi$ is a dummy flow variable.
@@
If the perturbation of a dummy variable ${\phi}^{\prime\prime}_{i}$ of a control volume can be represented as the harmonic form:
$\phi^{\prime\prime}_{i}=\sum^{l=N_{h}}_{l=1}(\hat{\phi}_{i,l}e^{I\omega_{i,l}t}+\hat{\phi}_{i,-l}e^{-I\omega_{i,l}t})$ (2)
@@
where $l$ is the harmonic index, $N_{h}$ is the number of harmonics, $I$ is the imaginary unit $\sqrt{-1}$, $\omega_{l}$ is the angular frequency of the $l^{\text{th}}$ harmonic and $t$ is time. Using the flow decomposition and the representation of the flow perturbation in the harmonic form, the coupled system of steady mean flow and harmonic flow perturbations in the frequency domain for a control volume can be obtained and solved with pseudo-time stepping. This can be represented symbolically as:
$\mathbf{M}\frac{\partial\mathbf{W}_{i}}{\partial t^{*}}+\mathbf{R}_{i}+\text{DF}_{i}$ $=0$ (3)
$\mathbf{M}\frac{\partial\mathbf{\hat{W}}_{l,i}}{\partial t^{*}}+\mathbf{\hat{R}}_{l,i}$ $=0$ (4)
@@
in which $t^{*}$ is the pseudo-time, $l$ is the harmonic index, $V$ is the volume of the current control volume and $\mathbf{M}$ is the Jacobian of the transformation from primitive variable $\mathbf{W}$ to conservative variables $\mathbf{Q}$. $\mathbf{R}$ is the non-linear residual of the mean flow and $\mathbf{\hat{R}}_{l}$ is the non-linear residual for the $l^{th}$ harmonic of the time-linearized flow. DF represents the contribution of deterministic fluxes [29], which takes into account the coupling between mean flow and harmonic flow perturbations. and can be determined using the mean flow variable and harmonic flow perturbations [3]. The nonlinear residual of the time-linearized flow $\mathbf{\hat{R}}_{l,i}$ of control volume $i$ is:
$\mathbf{\hat{R}}_{l,i}=I\omega_{l,i}\mathbf{M}V_{i}\mathbf{\hat{W}}_{l,i}+\frac{\partial\mathbf{R}}{\partial\mathbf{W}}\mathbf{\hat{W}}_{l,i}$ (5)
in which $\frac{\partial\mathbf{R}}{\partial\mathbf{W}}\mathbf{W}^{\prime\prime}$ represents the linearized flux that is accumulated on all the faces of the control volume of node $i$, here the indices (i.e. $j$) are ignored and it is written symbolically for simplicity. In Hydra, the linearized flux is computed by performing an automatic differentiation of the subroutines that are related to the nonlinear residuals.
It should be noted that, although Hydra solves the conservative form of the NS equation, it works with primitive variables. This is why the coupled system in Equation 3 and Equation 4 is written in the form of primitive variables. The formulation of FNLH is preferred in Hydra over the NLH method, because the NLH method is formulated with time-averaged conservative variables and requires an extra step to convert time-averaged conservative variables to time-averaged primitive variables [2], but this has been avoided in FNLH and will lead to significantly less modifications of the baseline code.
@@
Similarly to the baseline steady flow solver, FNLH uses pseudo-time stepping to solve the mean flow (Equation 3) and the time-linearized harmonic flow (Equation 4). Each pseudo-time step is called a non-linear iteration or outer iteration. In this non-linear iteration, the harmonic flow perturbations are evaluated in the first place, DF is then computed using the available flow perturbations in the harmonic form, and the mean flow is then updated with the inclusion of DF. This procedure is illustrated by Fig. 3. It is noted that the cross-coupling of the harmonics is ignored in the current work. This will be explored in our future work.
Figure 3: Flow chart of FNLH for each non-linear iteration.
In a multistage setting of FNLH, bladerow interfaces for the mean flow and the time-linearized harmonic flow are treated separately. For the mean flow, the original mixing plane formulation [30] is corrected by DF to take into account the unsteady effect [3]. For the time-linearized flow, a unified formulation has been developed to transmit disturbances in a multi-row environment [4]. Quasi-3D non-reflective treatment [31] are applied to the mean flow and the time-linearized flow.
Convergence Acceleration Procedure of FNLH with The Generalized RK scheme
@@
Hydra uses the Runge-Kutta scheme to advance the solution in pseudo-time. For the baseline steady flow solver, a generalized Runge-Kutta (RK) scheme has been developed for both explicit [24] and implicit [32] schemes, but this procedure is not available to the time-linearized harmonic flow solver. In this paper such a generalized RK scheme is extended for FNLH to solve the mean flow and the time-linearized harmonic flows together. This will lead to a unified formulation to accelerate the convergence of FNLH. In the following, the generalized RK scheme for the mean-flow solver is described in the first place, and this is then extended to solve the time-linearized harmonic flow.
Generalized RK Scheme for The Mean Flow
The generalized RK scheme in Hydra is based on the hybrid five-stage scheme RK (5,3) [22], where the non-linear viscous residuals are evaluated at stages 1, 3 and 5. It can be written as:
$\mathbf{Q}^{n,(0)}$ $=\mathbf{Q}^{n}$
$\mathbf{Q}^{n,(k)}$ $=\mathbf{Q}^{n,(0)}-\Gamma\alpha^{(k)}\mathbf{K}^{-1}(\mathbf{R}^{(k)}-\mathbf{R}^{*}),k=\{1,2\cdots 5\}$
$\mathbf{Q}^{n+1}$ $=\mathbf{Q}^{n,(5)}$ (6)
In Equation Generalized RK Scheme for The Mean Flow, without losing generality, the control volume index $i$ has been ignored for simplicity, $n$ is the current pseudo-time step index, $k$ is the RK stage index, $\mathbf{R}^{(k)}$ is the non-linear residual of the $k^{\text{th}}$ RK stage, $\mathbf{R}^{*}$ is the additional term related to multigrid operation, the formulation of which can be found in Moinier [18]. For each RK stage, a solution update needs to be computed to advance the solution. Such a solution update can be worked out in an explicit way, and this is the classic explicit RK approach. In addition, one can also invert the Jacobian matrix in an approximate way to obtain the update of the solution for each RK stage. In this case, since there is no multigrid operation involved, $\mathbf{R}^{*}$ in Equation Generalized RK Scheme for The Mean Flow is set to zero. The matrix inversion can be realized with the incomplete lower-upper decomposition with zero level of fill-in. This technique can be used as a preconditioner for Newton-Krylov methods, and it can also be used as a smoother to work out solution updates. The generalized RK scheme allows both explicit and implicit schemes to be written in a unified way, according to Misev [25] the formulations of $\Gamma$ and $\mathbf{K}$ for explicit and implicit approaches can be written as:
where $\mathbf{J}$ is the residual Jacobian matrix $\frac{\partial{\mathbf{R}}}{\partial\mathbf{Q}}$. The residual Jacobian is assembled using the information of the immediate neighbors with first-order spatial accuracy, and hence it is an approximation of the true Jacobian. $\mathbf{P}$ is the diagonal block of the Jacobian matrix $\mathbf{J}$. For the explicit scheme, $\Gamma$ is the the Courant–Friedrichs–Levy (CFL) number $\sigma_{CFL}$, in Hydra $\sigma_{CFL}=2$ is recommended and used by default in this work. It should be noted that for the implicit scheme, $\Gamma$ no longer represents the CFL number, but it is related to a relaxation factor $\varepsilon$. The detailed derivation can be found in Ref. [32, 22], a value of 0.6 is suggested in the previous work [32].
@@
To improve the convergence rate and stability, the non-linear residual of the $k^{\text{th}}$ stage $\mathbf{R}^{(k)}$ is decomposed into two components, namely the inviscid contribution $\mathbf{R}^{(k),inv}_{i}$ and the dissipative or viscous contribution $\mathbf{R}^{(k),vis}$:
$\mathbf{R}^{(k)}$ $=\mathbf{R}^{(k),inv}+\mathbf{R}^{(k),vis}$ (9)
@@
$\mathbf{R}^{(k),inv}$ $=\mathbf{R}^{inv}(\mathbf{Q}^{n,(k-1)})$ (10)
$\mathbf{R}^{(k),vis}$ $=\beta_{k}\mathbf{R}^{vis}(\mathbf{Q}^{n,(k-1)})+(1-\beta_{k})\mathbf{R}^{(k-1),vis}$ (11)
$\mathbf{R}^{inv}(\mathbf{Q}^{n,(k-1)})$ is the inviscid contribution using the flow variable $\mathbf{Q}^{n,(k-1)}$, $\mathbf{R}^{(k),vis}$ is the viscous contribution and is a mixture of the viscous flux using the flow variable $\mathbf{Q}^{n,(k-1)}_{i}$ and the viscous residual $\mathbf{R}^{(k-1),vis}$ from the previous RK stage. The coefficients $\alpha^{(k)}$ and $\beta^{(k)}$ are summarized in Table 1. These coefficients are optimized to damp high-frequency errors and lead to better convergence when used with multigrid [33].